OpenStudy (anonymous):

Please check my answer to this Question. I want to know if you got the same answer. following three problems involve the resistance of a ten meter long gold wire that is 1 millimeter in diameter. what is the wires cross sectional area? 0.31 cm^2 what is the unit for the resistance? ohm what is the number of the gold wires resistance?0.3

OpenStudy (espex):

What did you use to get your cross-sectional area?

OpenStudy (anonymous):

the diameter

OpenStudy (espex):

And the formula?

OpenStudy (anonymous):

\[A_{circle} = \pi r^2 = \frac{ \pi}{4}d^2\]

OpenStudy (anonymous):

no idea how you got the area that you did

OpenStudy (espex):

Which is where I was going, your number does not match.

OpenStudy (anonymous):

my book doesn't use that formula. it just gives me the formula for resistance